Admitting New Partner With Bonus

Cody Jenkins and Lacey Tanner formed a partnership to provide landscaping services. Jenkins and Tanner shared profits and losses equally. After all the tangible assets have been adjusted to current market prices, the capital accounts of Cody Jenkins and Lacey Tanner have balances of $32,000 and $42,000, respectively. Valeria Solano has expertise with using the computer to prepare landscape designs, cost estimates, and renderings. Jenkins and Tanner deem these skills useful; thus, Solano is admitted to the partnership at a 30% interest for a purchase price of $20,000.

a. Determine the recipient and amount of the partner bonus.
$  

b. Provide the journal entry to admit Solano into the partnership. For a compound transaction, if an amount box does not require an entry, leave it blank.

c. Why would a bonus be paid in this situation?

Apparently, Jenkins and Tanner value   offered by Solano.
Answer #1
Answer

a

Total capital after Solano's admission = 32000+42000+20000= $94,000
Solano's share = 94,000*30%= $28,200
Bonus = 28200-20000 = $8,200 Valeria Solano
b
Cash $     20,000
Cody Jenkins, Capital $       4,100
Lacey Tanner,Capital $       4,100
      Valeria Solano Capital $     28,200
( To record admission of admit Solano)
c
Apparently, Jenkins and Tanner value the expertise offered by Solano.

Please like